# Gravitational wave probes of particle dark matter: a review

### v3

@@ III.1.4 Stochastic summing method
Depending on the type of DM signal assumed to be interacting with the interferometers, $\bar{\lambda}_{\mathrm{up}}$ can be equated to the following physical normalized (by the noise) DM signal amplitudes to obtain the upper limit on the corresponding coupling constant:
$\bar{\lambda}_{\mathrm{axion}}$ $=\frac{T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}g_{\alpha\gamma}\sqrt{\frac{\rho_{\rm DM}}{m^{2}}}\frac{m}{2k},\quad$
$\bar{\lambda}_{\mathrm{time}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}}{mL}\sin^{2}\left(\frac{mL}{2}\right),\quad$
$\bar{\lambda}_{\mathrm{space}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}\bar{v}}{2{\sqrt{2}}},\quad$
$\bar{\lambda}_{\mathrm{charge}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{|(Q/M)_{e}-(Q/M)_{\mathrm{in}}|}{2Lm}.$ (80)
where $L$ is the length of the interferometers and $(Q/M)_{e},(Q/M)_{\mathrm{in}}$ are the charge-to-mass ratios of the end mirrors and input mirrors, respectively. The above approach works if $T_{\text{coh}}\neq T_{\text{obs}}$, however in practice, this does happen, since the data are typically broken in to smaller chunks with $T_{\text{FFT}}\sim T_{\text{coh}}$. The likelihood equation derived in Nakatsuka et al. (2023) is actually unstable in this region, and so a numerical approach is used to estimate sensitivity and derive upper limits as in Abac et al. (2024).
III.1.5 Distinguishing amongst dark-matter models
Only a couple of works Miller et al. (2022a); Manita et al. (2024) have attempted to find ways of determining whether a particular type of DM interacted with GW detectors. In Miller et al. (2022a), the authors employ the Wiener filter Wiener et al. (1964) to compute the cross power across different detector pairs in the follow-up stages of the analysis, using different $T_{\text{FFT}}$, to exploit the ways that types of DM will couple to different interferometer baselines. Not only was the Wiener filter shown to produce low residuals when the signal in the data matched the model, and high values when the data were filtered with the wrong signal but also it could, with high confidence, veto candidates returned in a real search of O3 data. Ref. Manita et al. (2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different. In both cases, cross power is important, since the individual power spectra are more or less indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
III.2 Ground-based interferometers
Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al. (2022c); Vermeulen et al. (2021); Göttel et al. (2024); Abac et al. (2024). We summarize the results that have arisen from each of the existing ground-based interferometers.
III.2.1 LIGO, Virgo and GEO600
In Abbott et al. (2022c), two methods were used to search for ultralight dark matter: cross-correlation and excess power. Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method. On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
In Fig. 3, we show the upper limits on the coupling constant $\epsilon^{2}$ for both methods, in comparison to existing experiments. We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments. These upper limits were derived for cross-correlation and excess power by following the procedure outlined in Section III.1.1 and Section III.1.2, respectively. Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al. (2021).
Figure 3: Adapted from Abbott et al. (2022c). Upper limits obtained from analyzing LIGO O3 data on the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors. MICROSCOPE Bergé et al. (2018) and Eot-wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al. (2008). To produce limits on dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
Additionally, a recent search for scalar bosons interacting with LIGO was performed Göttel et al. (2024) using LPSD, resulting in tight constraints on the coupling constants given in Eq. 27. In Fig. 4, we can see, in blue, the constraints that have been derived, which supersede, at least a low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al. (2021). The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search. These upper limits were derived for by following the procedure outlined in Section III.1.3 and Section III.1.2.
Figure 4: Taken from Göttel et al. (2024). 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors. a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively. The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al. (2021); Aiello et al. (2022); Aharony et al. (2021); Savalle et al. (2021); Kennedy et al. (2020); Antypas et al. (2019, 2021); Tretiak et al. (2022); Oswald et al. (2022); Campbell et al. (2021); Beloy et al. (2021); Zhang et al. (2023); Fukusumi et al. (2023); Sherrill et al. (2023) are shown in thin grey, and constraints from searches for “fifth forces” Bergé et al. (2022); Hees et al. (2018) are depicted by the dashed red lines.
III.2.2 KAGRA
KAGRA is another interferometer in Japan that, while not yet as sensitive to GWs as LIGO and Virgo, can be used to search for DM. In particular, the different materials used in the mirrors in KAGRA can be exploited to be sensitive to vector DM that couples to baryon-lepton number Michimura et al. (2020), since the signal-to-noise ratio is enhanced for objects with different charge-to-mass ratios, as indicated by Eq. 80. Furthermore, this kind of DM can actually be seen in non-GW channels, such as the auxiliary length channels PRCL and MICH. The raw outputs of PRCL and MICH are time-varying displacement sof the forms (we also write the GW channel for comparison, $\delta L_{\rm DARM}$) Abac et al. (2024):
$\delta L_{\rm DARM}$ $=$ $\delta(L_{\rm x}-L_{\rm y}),$ (81)
$\delta L_{\rm MICH}$ $=$ $\delta(l_{\rm x}-l_{\rm y}),$ (82)
$\delta L_{\rm PRCL}$ $=$ $\delta[(l_{\rm x}+l_{\rm y})/2+l_{\rm p}],$ (83)
In KAGRA, $L_{x}=L_{y}=3000$m is the arm length, $l_{x}=26.7$m is the distance between the beam splitter and the input $x$ mirror, $l_{y}=23.3$m is the distance between the beam splitter and the input $y$ mirror, and $l_{p}=41.6$m is the the length of the power recycling cavity. All of these lengths are much shorter than the DM coherence length $L_{\text{coh}}$. See Fig. 1 of Abac et al. (2024) for a schematic of the interferometer.
These two channels contain both fused silica mirrors used for the beam splitter and power recycling mirrors, and cryogenic sapphire test masses. The charge-to-mass ratios for baryon-lepton number differ by $\sim 9\times 10^{-3}$ for fused silica and sapphire, while for baryon number alone, these quantities differ by only $\sim 4\times 10^{-5}$, showing a large enhancement of the baryon-lepton DM interaction signal.
No significant signals were found in this analysis of KAGRA O3 data Abac et al. (2024), so upper limits, as shown in Fig. 5, were calculated following the procedure outlined in Section III.1.4. The upper limits are shown for each channel analyzed, and many narrow-band noise artifacts are visible, which tends to limit the sensitivity of this search at low freqeuncies. The constraints are not as strong as those that come from LIGO; however, future searches with KAGRA are expected to yield powerful constraints on this type of DM due to the difference in charge-to-mass ratios of the mirrors Michimura et al. (2020).
Figure 5: Taken from Abac et al. (2024). 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line). Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced, and is more visible in these particular channels than in the ordinary strain one. In the low-mass (low-frequency) range, many narrow-band noise disturbances are seen, which arise due to unknown sources.
III.3 Space-based interferometers
Spaced-based GW interferometers, such as LISA Danzmann and Team (2003); Babak et al. (2021), Taiji Hu and Wu (2017), Tianqin Luo et al. (2016) and DECIGO Hu and Wu (2017), will hopefully fly within the next 10-15 years. The exquisite sensitivity of GWs in the $\mu$Hz to mHz band also allows the prospect to probe ultralight dark matter with masses of [$10^{-19},10^{-14}$] eV, a few orders of magnitude lower than those currently searched for in ground-based interferometers Kim (2023). Until now, however, only a protoype for LISA, called LISA Pathfinder, has been flown as a “proof-of-concept” mission Armano et al. (2016); Vetrugno (2017); Armano et al. (2018). Nevertheless, pilot searches can be performed on LISA Pathfinder data as a way to prepare for when space-based detectors fly, which allows for the development of robust data analysis pipelines and for handling of the peculiarities of future data, e.g. glitches, downtimes, etc.
@@ III.3 Space-based interferometers
Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to $\mathcal{O}(10^{6})$ km), and only one arm existed. However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
@@
After this search was performed, however, it was realized that the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously, which was the relative acceleration of the two test masses Frerick et al. (2024). A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig. 6, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses. We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled .(decoherence). The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
@@
Figure 6: Taken from Frerick et al. (2024). Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder. The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” . Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials. Constraints from other experiments and future projections for LISA are shown as well.
@@
Only B-L constraints are shown from Frerick et al. (2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than provided by MICROSCOPE. However, many conservative assumptions were made in Frerick et al. (2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
III.4 Pulsar timing arrays
Because pulsars are extremely accurate clocks, they can also be used to search for GWs, which will distort the spacetime between earth and pulsars. Practically speaking, GWs will cause a change in the times of arrivals of pulses with respect to that in the absence of GWs. Of course, other effects between the pulsars and earth must be taken into account, e.g. interstellar medium dispersion and scattering, instabilities in the rotation of the pulsar itself, which makes these measurements particularly challenging. Nevertheless, pulsar timing arrays can be formed by cross correlating the times of arrivals from multiple pulsars in the galaxy, mitigating many of these effects.
Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers. It can also induce stochastic perturbations in the metric, which would also induce time-varying changes in the times of arrivals of pulses Kim and Mitridate (2024), but generally could be seen in interferometers with long baselines Kim (2023).
In Section II, we discussed various models of scalar DM that can be constrained with pulsar timing arrays. We again divide this section in the different types of DM that can be probed with pulsar timing arrays.
III.4.1 Gravitationally-interacting ultralight dark matter
@@
As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs. Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction. In Fig. 8, we show recent constraints on ultralight scalar DM in two ways: (1) the minimum detectable strain amplitude as a function of DM mass, and (2) the fraction of DM that ultralight dark matter could compose. We can see that the masses below $\sim 10^{-23.2}$ eV can be well constrained by pulsar timing arrays, and that ultralight dark matter cannot make up all of DM at these masses. Fig. 8 presents upper limits in three scenarios: (1) uncorrelated, (2) correlated and (3) pulsar-correlated.
@@
    “Correlated” means that the earth-pulsar distance and the average distance between pulsars is smaller than $L_{\text{coh}}$, and that the DM $L_{\text{coh}}$ comprises the inner 20 kpc region of the galaxy for which galaxy rotation curves have been used to test the DM hypothesis. This means that each pulsar experiences the same phase of the ultralight dark-matter signal. In this case, the same coherence patch and local DM energy density are measured.
@@
    “Pulsar-correlated” means that $L_{\text{coh}}$ is smaller than the radius of the galactic center sampled by experiments that derive galaxy rotation curves. However, estimates of the DM energy density are from different coherence patches and thus measurements averaged over different patches.
@@
Figure 8: Taken from Smarra et al. (2023). 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the times of arrivals of pulses in an analogous way as GWs do. The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency. $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$. For comparison, upper limits on these quantities are shown from previous searches Porayko and Postnov (2014); Porayko et al. (2018). In the right panel, we zoom in on the excluded ultralight dark matter masses. The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density . Scalar ultralight dark-matter particles with mass $-24.0<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40$ of $\rho_{\text{DM}}$, while particles with masses $-23.7<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ of $\rho_{\text{DM}}$.
Additionally, NANOGrav data can be used to constrain this kind of ultralight dark matter. In Fig. 9, a similar constraint is shown on a quantity that relates to the fraction of the DM energy density that gravitationally-interacting DM could compose. $S_{\phi}$ is the spin of DM, which can be 0 or 1 in this model. NANOGrav writes the timing residuals for the $I^{\rm th}$ pulsar as:
$h_{I}(t)$ $=\sum_{i}A_{E,I}^{i}(\bm{x}_{E})\sin(\omega t+\gamma_{E}^{i})$ (84)
$+A_{P,I}^{i}(\bm{x}_{P,I})\sin(\omega t+\gamma_{P,I}^{i})\,,$ (85)
where $i$ indicates the polarization ($=0$ for spin-0 and $=1,2,3$ for spin-1). This expression is divided into “earth” and “pulsar” terms with amplitudes and arbitrary phases of $A_{E,I}^{i}(\bm{x}_{E})$ and phase $\gamma_{E}^{i}$, and $A_{P,I}^{i}(\bm{x}_{P,I})$ and time-independent phase $\gamma_{P,I}^{i}$, respectively. These phases depend on the positions of the earth and the pulsar with respect to the ultralight dark-matter field (see Eq. 60 and 61 of Afzal et al. (2023)).
Like the EPTA, NANOGrav assumes a model of ultralight dark matter that induces metric perturbations and therefore changes in the times of arrivals of pulses Khmelnitsky and Rubakov (2014). The amplitude of the signal that is constrained is similar to Eq. 64:
$A_{E,I}^{i}(\bm{x}_{E})=A_{P,I}^{i}(\bm{x}_{E})=A_{\rm grav}=\left(2S_{\phi}+1\right)\displaystyle\frac{\pi G\rho_{\text{DM}}}{2m_{\rm DM}^{2}}\hat{\phi}^{2}(\bm{x})$ (86)
where $S_{\phi}=0$ or 1 for scalar and vector DM, respectively.
@@
Fig. 9 shows the NANOGrav constraints on this amplitude, which can be compared to the right-hand panel of Fig. 8. NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
Figure 9: Taken from Afzal et al. (2023). Upper limits on the local ULDM density for the correlated (solid line) and uncorrelated (dashed line) signals at the $95\%$ credible level using data from the NANOGrav pulsar timing array. The model of DM considered here is the same as that in Fig. 8. The gray dashed line indicates the predicted DM abundance.
III.4.2 Variation of fundamental constants constraints
From the most recent NANOGrav dataset, constraints on ultralight dark matter that causes fundamental constants to oscillate Afzal et al. (2023) can be set as well for masses below $10^{-22}$ eV. In the case that ultralight dark matter causes changes in the moment of inertia of the pulsar, Fig. 10 depicts constraints on the couplings of the scalar DM particle to up- and down- quarks weighted by their masses ($\hat{m})$, electrons $(d_{e})$, muons $(d_{\mu})$, photons $(d_{\gamma})$ and gluons $(d_{g})$. To derive these constraints, the simplest model of the equation-of-state of the neutron star is assumed, as outlined in Section II.3.2, and, when constraining one particular coupling constant, the rest are assumed to be zero.

### v4

@@ III.2.3 LISA Pathfinder
Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed. However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
After this search was performed, however, it was realized that data from a different channel would produce tighter constraints, particularly in the case of baryon-lepton coupling. Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick et al. (2024). This echoes the logic presented in Section III.2.2 to use KAGRA to search for dark photons coupling to baryon-lepton number.
@@
A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig. 7, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses. We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled (decoherence). The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
@@
Figure 7: Taken from Frerick et al. (2024). Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder (dark blue region). The physics behind this form of DM has been discussed in Section II.4, and constrains the same model as in Figs. 6 and 4, though at much lower boson masses. The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” (dark red region). Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials. Constraints from other experiments (Eöt-Wash and MICROSCOPE) and future projections for LISA, MAGIS-100 and torision balance are shown as well. Details on the curves coming from STE-QUEST and Asteroids can be found in O’HARE (2020).
@@
Only B-L constraints are shown from Frerick et al. (2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE. However, many conservative assumptions were made in Frerick et al. (2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
III.3 Pulsar timing arrays
Because pulsars are extremely stable astrophysical clocks, they can be used to search for GWs, which perturb the spacetime between Earth and pulsars. In practical terms, GWs induce fluctuations in the ToAs of radio pulses, relative to what would be expected in their absence. However, interpreting these ToA shifts is complicated by other astrophysical and instrumental effects – such as dispersion and scattering in the interstellar medium, and intrinsic rotational instabilities of the pulsars themselves – which makes precision measurements challenging. Nevertheless, these effects can be mitigated by cross-correlating data from many pulsars in the sky to form a pulsar timing array.
As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky and Rubakov (2014), though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers. Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim and Mitridate (2024), which could lead to further timing variations over long baselines Kim (2023).
Throughout this section, we will see that, independently of the physical mechanism of coupling of DM to gravity or the standard model, the shapes of the constraints on the coupling constant appear similar. This is because in these cases, we are looking for an almost sinusoidal signal embedded in pulsar timing array data with particular noise characteristics that are independent of the DM model. Though the underlying physics behind each DM model is different, the observable – shifts in ToAs– is identical, though with different amplitudes that relate to the coupling constant of DM. This connects to our previous comments in the introductory paragraphs of Section III: the signal arising from DM coupling to gravity or the standard model is the same – a quasi-sinusoidal oscillation of spacetime, of materials, of test masses, etc. – but the cause and therefore the interpretation of time-varying ToA in pulsar timing arrays or time-varying displacements of test masses in ground- or space-based GW interferometers is fundamentally different, which leads to different constraint plots whose curves look similar.
In Section II.3, we discussed the theoretical mechanisms by which scalar DM could leave signatures on pulsar timing array data, and in Section III.2 we described how related effects could be searched for with ground-based interferometers. In contrast to that section, here we organize existing constraints from pulsar timing arrays according to the specific physical mechanism by which scalar DM couples to the standard model or to gravity. Each mechanism leads to a distinct signature in the pulsar timing residuals. We treat these mechanisms independently: if we put constraints on DM coupling to dilatons in the context of one mechanism, we assume that all other mechanisms do not contribute
III.3.1 Gravitationally-interacting ultralight dark matter
@@
As discussed in Section II.6, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs. Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction. In Fig. 8, we show recent constraints on ultralight scalar DM in two ways: (1) the minimum detectable strain amplitude as a function of DM mass, and (2) the fraction of DM that ultralight dark matter could compose. We can see that the masses below $\sim 10^{-23.2}$ eV can be well constrained by pulsar timing arrays, and that ultralight dark matter cannot make up all of DM at these masses. Fig. 8 presents upper limits in three scenarios: (1) uncorrelated, (2) correlated and (3) pulsar-correlated.
@@
    “Correlated” means that the earth-pulsar distance and the average distance between pulsars is smaller than $L_{\text{coh}}$, and that the DM $L_{\text{coh}}$ comprises the inner 20 kpc region of the galaxy for which galaxy rotation curves have been used to test the DM hypothesis. This means that each pulsar experiences the same phase of the ultralight dark-matter signal. In this case, the same coherence patch and local DM energy density are measured: the amplitude of the ultralight dark-matter interaction signal can be attributed completely to measurement of $\rho_{\text{DM}}$ coming from galaxy rotation curves [^6] This case is directly analogous to measurements of ultralight dark-matter couplings in ground- and space-based GW interferometers, discussed in Section III.2, where the arm lengths are negligible compared to $L_{\text{coh}}$, so all test masses experience the same ultralight dark-matter phase..
@@
    “Pulsar-correlated” means that $L_{\text{coh}}$ is larger than the earth-pulsar and inter-pulsar separations, so $\hat{\phi}_{P}^{2}=\hat{\phi}_{E}^{2}$, but smaller than the galactocentric radius sampled by galaxy rotation curves. In this case, pulsars and the Earth probe the same coherence patch, but the ultralight dark-matter amplitude measured with pulsars cannot be identified one-to-one with the DM energy density inferred from galaxy rotation curves, since the latter averages over many independent coherence patches.
@@
Figure 8: Taken from Smarra et al. (2023). 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the ToAs of pulses in an analogous way that GWs do. The physics behind this form of DM has been discussed in Section II.6. The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency. $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$. For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko and Postnov (2014); Porayko et al. (2018). Furthermore, the dashed line indicates the signal amplitude assuming that this form of DM comprises all of DM; thus, curves that lie above this dashed line do not represent physical constraints. Upper limits on the amplitude are shown for the three cases outlined in Section III.3.1. In the right panel, we zoom in on the excluded ultralight dark matter masses. The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density. The right panel shows that scalar ultralight dark-matter particles with mass $24.0<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40\%$ of $\rho_{\text{DM}}$, while particles with masses $23.7<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ % of $\rho_{\text{DM}}$.
@@
Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter Afzal et al. (2023), though the constraints are weaker than from the EPTA. In particular, NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
III.3.2 Constraints on scalar conformal dark matter
While the previous subsection considered generic strain-like gravitational interactions of scalar ultralight dark matter, scalar conformal couplings represent a distinct class of models in which ULDM interacts through the trace of the energy–momentum tensor, As discussed in Section II.6. This model of DM can also be constrained with pulsar timing array data, and in particular Ref. Smarra et al. (2024) uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
In Fig. 9, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data. However, despite showing upper limits (that assume no signal), Ref. Smarra et al. (2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses. It is worth noting that the effective change in the neutron-star moment of inertia in this scenario arises from the universal rescaling of the metric due to the conformal coupling, and is therefore distinct from the composition-dependent moment of inertia variations discussed in Section II.3, where scalar fields couple directly to quarks and nucleons. The constraints arising from the interaction of dilatons with standard-model particles in neutron stars will be discussed in the next subsection.
Figure 9: Taken from Smarra et al. (2024). Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al. (2024) to estimate $\alpha$ (brown dashed lines). The physics behind this form of DM has been discussed in Section II.6. $\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$. Here, this model of DM would change the moment of inertia of neutron stars used in pulsar timing arrays, resulting in changes of the ToA of the pulses. If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in Smarra et al. (2023). AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal et al. (1998).
III.3.3 Variation of fundamental constants constraints
[…]